# Well-posedness of Hibler’s parabolic-hyperbolic sea ice model

### v2

@@ 1. Introduction
It is the aim of the present paper to prove the local strong well-posedness of the parabolic-hyperbolic variant of Hibler’s model, where the most commonly used regularization for the stress tensor is employed. Unlike the models under consideration in [6] or [23], this is precisely the model typically studied in numerical simulations.
@@ 2. Hibler’s sea ice model and main result
For the ratio $e>1$ of the major to minor axes of the elliptical yield curve of the principle components of the stress, we introduce
@@
The resulting bulk and shear viscosities then take the shape
@@
for densities $\rho_{\mathrm{atm}}>0$ and $\rho_{\mathrm{ocn}}>0$, drag coefficients $C_{\mathrm{atm}}>0$ and $C_{\mathrm{ocn}}>0$, velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ of the atmosphere and the ocean and rotation matrices $R_{\mathrm{atm}}$ and $R_{\mathrm{ocn}}$.
@@
where $q\in(1,\infty)$. With regard to the initial values, for $p\in(1,\infty)$, we then introduce the time trace space $\mathrm{X}_{\gamma}=(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$ taking the shape
@@
In the above, the subscript D indicates Dirichlet boundary conditions. If $p$, $q\in(1,\infty)$ satisfy
(2.6) $\frac{1}{p}+\frac{1}{q}<\frac{1}{2},$
then it follows from Sobolev embeddings, see for example [35, Theorem 4.6.1], that
@@
Let $q\in(1,\infty)$, and let $T_{1}>0$ sufficiently large to be determined precisely later. We make the following assumptions on the data.
@@
    The wind velocity $V_{\mathrm{atm}}$ and the ocean velocity $V_{\mathrm{ocn}}$ satisfy $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}\in\mathrm{L}^{\infty}(0,T_{1};\mathrm{L}^{2q}(\Omega)^{2})$.
@@
    For the sea surface dynamic height $H$, it is valid that $\nabla_{\mathrm{H}}H\in\mathrm{L}^{\infty}(0,T_{1};\mathrm{L}^{q}(\Omega)^{2})$.
@@ 3. Transformation to Lagrangian coordinates
Besides, we compute
@@
Next, we study the continuous dependence of the solution on the initial data.
@@ 4. Linear theory
In this section, we address the linear theory. More precisely, we establish the boundedness of the $\mathcal{H}^{\infty}$-calculus of the operator matrix associated to the linearization of the parabolic-hyperbolic sea ice model. This result is of independent interest. As a corollary, we especially obtain the maximal $\mathrm{L}^{p}$-regularity for the fixed point argument.
@@
In the sequel, we fix $\omega>\omega_{0}$ for $\omega_{0}>0$ from Section 4. The maximal regularity constant $C_{\mathrm{MR}}>0$ can in particular be chosen independent of $T>0$ provided homogeneous initial values are considered. In order reduce the considerations to this situation, we now discuss the so-called reference solution capturing the initial values. More precisely, for $u_{0}\in V$, the resulting system reads as
@@ 6. Concluding remarks
The bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix $A(u_{1})$ from (4.3) as proved in Theorem 4.1 paves the way for interesting applications such as the theory of Banach scales due to Amann [2, Chapter 5] or stochastic maximal regularity, see also [27] for more details on this property.
@@
Acknowledgements The author would like to thank Tim Binz, Karoline Disser and Matthias Hieber for fruitful discussions. Moreover, he acknowledges the support by DFG project FOR 5528. Most of the results in this paper are contained in modified form in Chapter 6 of the PhD thesis [5] of the author.

### v3

@@ 1. Introduction
It is the aim of the present paper to prove the local strong well-posedness of the parabolic-hyperbolic variant of Hibler’s model, where the most commonly used regularization for the stress tensor is employed. Unlike the models considered in [6] or [23], this is the model typically studied in numerical simulations.
@@
For convenience of the reader, we briefly elaborate on the function spaces used in this paper. Consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with boundary of class $\mathrm{C}^{2}$, and suppose that $G$ either represents $\Omega$, or an interval of the form $(0,T)$, for $T>0$. With $p\in[1,\infty]$ and $k\in\mathbb{N}$, we denote by $\mathrm{L}^{p}(G)$ and $\mathrm{W}^{k,p}(G)$ the Lebesgue and Sobolev spaces on $G$, respectively. The space $\mathrm{W}_{0}^{1,q}(\Omega)$, $q\in(1,\infty)$, contains the functions in $\mathrm{W}^{1,q}(\Omega)$ whose trace vanishes on $\partial\Omega$. Moreover, for $l\in\{0,1\}$ and $\alpha\in(0,1)$, the space $\mathrm{C}^{l,\alpha}(\overline{\Omega})$ consists of all functions $f\in\mathrm{C}^{l}(\overline{\Omega})$ such that all partial derivatives $\partial^{\beta}f$ of orders $|\beta|\leq l$ are $\alpha$-Hölder continuous on $\overline{\Omega}$. By $\mathrm{C}_{\mathrm{b}}^{1}([0,\infty))$, we denote the functions in $\mathrm{C}^{1}([0,\infty))$ that are bounded on $[0,\infty)$, and whose derivative is additionally bounded on $[0,\infty)$. The space $\mathrm{BUC}([0,T])$ represents the bounded and uniformly continuous functions on $[0,T]$. For $s>0$ and $p$, $q\in(1,\infty)$, we use $\mathrm{B}_{qp}^{s}(\Omega)$ for the Besov spaces on $\Omega$. Note that they can be obtained from the Lebesgue and Sobolev spaces by real interpolation, see e. g. [35, Sec. 2.4 and 4.3.1]. More precisely, for $s$, $p$, $q$ as above, we have $\mathrm{B}_{qp}^{s}(\Omega)=(\mathrm{L}^{q}(\Omega),\mathrm{W}^{2,q}(\Omega))_{\frac{s}{2},p}$. Finally, for a Banach space $\mathrm{X}$, we denote by $\mathrm{L}^{p}(0,T;\mathrm{X})$ and $\mathrm{W}^{1,p}(0,T;\mathrm{X})$ the respective Bochner spaces.
@@ 2. Hibler’s sea ice model and main result
It can be found in [13, p. 822, eq. (17)]. For the ratio $e>1$ of the major to minor axes of the elliptical yield curve of the principle components of the stress, we introduce
@@
The bulk and shear viscosities are as made precise in [13, p. 819, eq. (7) and (8)], and they take the shape
@@
for constant densities $\rho_{\mathrm{atm}}>0$ and $\rho_{\mathrm{ocn}}>0$, constant drag coefficients $C_{\mathrm{atm}}>0$ and $C_{\mathrm{ocn}}>0$, velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ of the atmosphere and the ocean and rotation matrices $R_{\mathrm{atm}}$ and $R_{\mathrm{ocn}}$.
@@
where $q\in(1,\infty)$. In the sequel, we will assume that $p$, $q\in(1,\infty)$ satisfy
(2.5) $\frac{1}{p}+\frac{1}{q}<\frac{1}{2}.$
With regard to the initial data, for $p$, $q\in(1,\infty)$ fulfilling (2.5), we then introduce the time trace space $\mathrm{X}_{\gamma}=(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$ taking the shape
@@
In the above, the subscript D indicates Dirichlet boundary conditions. For $p$, $q\in(1,\infty)$ such that (2.5) holds true, it follows from Sobolev embeddings, see for example [35, Theorem 4.6.1], that
@@
Let $q\in(1,\infty)$. We make the following assumptions on the data.
@@
    The wind and ocean velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ satisfy $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}\in\mathrm{L}^{\infty}(0,t;\mathrm{L}^{2q}(\Omega)^{2})$ for all $t>0$.
@@
    For the sea surface dynamic height $H$, it is valid that $\nabla_{\mathrm{H}}H\in\mathrm{L}^{\infty}(0,t;\mathrm{L}^{q}(\Omega)^{2})$ for all $t>0$.
@@
The above assumptions on the integrability and regularity of the data are relatively mild, and it seems that they are physically reasonable. For examples of ice growth rate functions $f_{\mathrm{gr}}$, which generally depend on the season, we refer for instance to the ones considered by Hibler [13, p. 821].
@@ 3. Transformation to Lagrangian coordinates
Let us observe that the coefficients $a_{ij}^{klm}$ depend on the time $t$ via $\partial_{k}Y_{m}$. For simplicity of notation, we do not write this dependence explicitly. Besides, we compute
@@
Next, we study the continuous dependence of the solution on the initial data. As a by-product, we also obtain the local uniformity of the time of existence $T>0$ with respect to the initial data.
@@ 4. Linear theory
In this section, we address the linear theory. More precisely, we establish the boundedness of the $\mathcal{H}^{\infty}$-calculus of the operator matrix associated to the linearization of the parabolic-hyperbolic sea ice model. This result is of independent interest due to its interesting functional analytic consequences for the operator matrix such as a characterization of its fractional power domains, or stochastic maximal regularity. We refer here also to the discussion in Section 6. As a corollary, we especially obtain the maximal $\mathrm{L}^{p}$-regularity. Note that this (weaker) property is sufficient for the fixed point argument.
@@
In the sequel, we fix $\omega>\omega_{0}$ for $\omega_{0}>0$ from Section 4. Note that the maximal regularity constant $C_{\mathrm{MR}}>0$ from Section 4 generally depends on $T$. However, it can be chosen independent of $T>0$ provided homogeneous initial values are considered. In order reduce the considerations to this situation, we now discuss the so-called reference solution capturing the initial values. More precisely, for $u_{0}\in V$, the resulting system reads as
@@ 6. Concluding remarks
The bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix $A(u_{1})$ from (4.3) as proved in Theorem 4.1 paves the way for interesting applications such as the theory of Banach scales due to Amann [2, Chapter 5] or the consideration of the problem in the stochastic setting via stochastic maximal regularity, see also [27] for more details on this property.
@@
Acknowledgements. The author would like to thank Tim Binz, Karoline Disser and Matthias Hieber for fruitful discussions. Moreover, he acknowledges the support by the German National Academy of Sciences Leopoldina through the Leopoldina Fellowship Program with grant number LPDS 2024-07 and the support by DFG project FOR 5528. Most of the results in this paper are contained in modified form in Chapter 6 of the PhD thesis [5] of the author.
Data Availability Statement. The manuscript has no associated data.